# Align & Invert: Solving Inverse Problems with Diffusion and Flow-based Models via Representational Alignment

### v1

@@ 4 Theoretical Results
Recent advances in generative modeling have demonstrated that objectives defined in deep feature spaces, rather than purely pixel-based discrepancies, produce reconstructions that better capture perceptual similarity (dosovitskiy2016generating; zhang2018unreasonable; fu2023dreamsim). Motivated by this observation, our goal in this section is to clarify how the Repa regularizer fits within this broader class of feature-based objectives. In particular, we present two theoretical results that characterize how Repa acts on (i) the DINOv2-space representations of reconstructed images and (ii) the hidden representations of the diffusion model itself.
First, we show that the Repa term can be viewed as an empirical surrogate for minimizing a feature-space divergence specifically, the maximum mean discrepancy (MMD) between the DINOv2 features of reconstructed images and those of the ground truth. Second, we study how Repa affects the hidden states of the diffusion model during the reverse process. Under suitable assumptions, we show that a Repa update moves the current diffusion representation toward the representation that the model would produce for the clean image. This provides a complementary view: Repa influences not only the DINOv2 representation of the reconstruction but also the internal representations that guide the sampling trajectory.
@@
Definition 1 (Misalignment between DINOv2 and diffusion representations).
@@
$\mathrm{MisREPA}(x,t):=\frac{1}{N}\sum_{n=1}^{N}\bigl\|f_{\mathrm{DINOv2}}^{[n]}(x)-g_{\phi}\!\bigl(\mathrm{DiffEnc}^{[n]}(x,t)\bigr)\bigr\|_{2}^{2},$
where $f_{\mathrm{DINOv2}}^{[n]}(x)$ and $\mathrm{DiffEnc}^{[n]}(x,t)$ denote the $n$-th patch features from the DINOv2 encoder and the diffusion model encoder, respectively. The mapping $g_{\phi}$ is the projection head introduced in Section 2.2.
Remark.
When the timestep argument is omitted, all $t$-dependent quantities are evaluated at $t=0$. In particular, $\mathrm{MisREPA}(x)\equiv\mathrm{MisREPA}(x,0)$ and $\mathrm{DiffEnc}^{[n]}(x)\equiv\mathrm{DiffEnc}^{[n]}(x,0)$.
Definition 2 (Proxy approximation error).
For a ground-truth image $x$ and its proxy approximation $\bar{x}$ (defined in Section 3), define the proxy approximation error as
$\mathrm{ApproxErr}(x,\bar{x}):=\frac{1}{N}\sum_{n=1}^{N}\bigl\|f_{\mathrm{DINOv2}}^{[n]}(x)-f_{\mathrm{DINOv2}}^{[n]}(\bar{x})\bigr\|_{2}^{2}.$
In practice, both quantities defined above are typically small. The feature misalignment $\mathrm{MisREPA}(x,t)$ is mitigated through the explicit alignment of DINOv2 and diffusion encoder representations during training (yu2024repa), while the approximation error $\mathrm{ApproxErr}(x,\bar{x})$ is reduced by constructing proxy features $\bar{x}$ that serves as reliable substitutes for the unavailable ground-truth features (see Section 3). Thanks to the robustness of DINOv2 features to common image degradations (oquab2023dinov2), the resulting approximation gap remains small across the inverse problems we study.
Feature-Space Alignment: Repa as an MMD Surrogate
Building on the definitions introduced above, we now establish a connection between the Repa regularizer and the statistical discrepancy between DINOv2 features of clean and reconstructed images. The proposition that follows formalizes this link by expressing the expected Repa score in terms of a feature-space divergence.
Proposition 1.
Assume $x\sim p_{X}$ is a ground-truth image and $\hat{x}\sim p_{\hat{X}\mid Y}$ is its corresponding reconstruction, where $p_{\hat{X}\mid Y}$ is the distribution induced by any reconstruction method. Assume further that all DINOv2 feature embeddings are $\ell_{2}$–normalized. Let $\bar{x}$ be a proxy approximation of $x$ (Section 3), with $\bar{x}\sim p_{\bar{X}}$. Define the Repa alignment as
@@
and the empirical mean DINOv2 feature embedding
$\mu_{f}(x):=\frac{1}{N}\sum_{n=1}^{N}f_{\mathrm{DINOv2}}^{[n]}(x).$
Then the expected Repa alignment satisfies
@@
The proof is provided in the Appendix A.1. Proposition 1 establishes a connection between the representation alignment term and the maximum mean discrepancy (MMD) (arbel2019maximum), which quantifies perceptual divergence between feature distributions. This relationship indicates that maximizing the Repa term encourages the minimization of a divergence measure captured by the MMD together with additional penalties accounting for feature misalignment between diffusion and DINOv2 representations, as well as an approximation error arising from the use of proxy rather than ground-truth features for alignment.
@@
When these terms are sufficiently small, Proposition 1 implies that maximizing the Repa term effectively corresponds to minimizing the maximum mean discrepancy (MMD) between the DINOv2 feature distributions of reconstructed and ground-truth images. This interpretation reveals that Repa regularization offers a principled, representation-based mechanism for enforcing perceptual consistency, by guiding generated images to align with the semantic structure captured by the DINOv2 encoder.
Effect of REPA on Internal Diffusion Representations
Having characterized the feature-space effect of Repa, we now turn to its influence on the representations of the diffusion model itself. The next result formalizes how a single Repa update impacts the model’s internal representation relative to that of the clean image.
Proposition 2.
Consider an inverse problem in which an observation $y\in\mathbb{R}^{m}$ is generated from a ground-truth image $x\sim p_{X}$, and let $\bar{x}\sim p_{\bar{X}}$ be the proxy approximation used during REPA (Section 3). Let $z_{t}\in\mathbb{R}^{d}$ be the diffusion state at timestep $t$ produced by Algorithm 1 given $y$, and let $z_{t}^{\mathrm{(REPA)}}$ denote the diffusion state after applying the Repa regularization update. Let $z^{*}$ denote the latent variable associated with the clean image $x$, and define
$h_{t}=G_{2}(z_{t},t),\qquad h_{t}^{\mathrm{(REPA)}}=G_{2}(z_{t}^{\mathrm{(REPA)}},t),\qquad h^{*}=G_{2}(z^{*},0),$
where $G_{2}$ denotes the feature extractor returning the patch-level representation in layer $\ell$ of the model. Assume that the projection head $g_{\phi}$ is linear and that all DINOv2 patch embeddings are $\ell_{2}$–normalized. Then there exists a regularization weight $\lambda>0$ (Algorithm 1) such that the REPA update induces an approximate contraction, i.e., there exists a constant $C_{1}\in(0,1)$ and a constant $C_{2}\geq 0$ satisfying
$\|h_{t}^{\mathrm{(REPA)}}-h^{*}\|_{F}\;\leq\;C_{1}\,\|h_{t}-h^{*}\|_{F}+C_{2}\!\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right).$
The proof is provided in Appendix A.1. Proposition 2 characterizes how the Repa update modifies the hidden representations of the diffusion model. The inequality shows that a single Repa step contracts the distance between the current representation $h_{t}$ and the clean-image representation $h^{*}$, up to two additive terms: the proxy error $\mathrm{ApproxErr}(x,\bar{x})$ and the encoder misalignment $\mathrm{MisREPA}(x)$. As explained, both quantities are expected to be small in practice. When this is the case, the result implies that Repa acts as a contraction on the model’s internal representations, steering them toward those associated with the clean image.
Intuitively, Proposition 1 shows that the Repa regularizer induces a reduction in the discrepancy between reconstructed and ground-truth images in the DINOv2 feature space an embedding space known to correlate well with human perceptual similarity (fu2023dreamsim). Proposition 2 complements this view by establishing that the REPA update also reduces the gap between the model’s internal hidden representation and that of the clean image.
To further illustrate these insights, Appendix A.5 compares Repa to a variant that enforces consistency directly in the DINOv2 feature space, without aligning the diffusion model’s internal states. The results show that Repa leads to higher perceptual quality, highlighting the benefit of guiding the sampling process through the model’s intermediate representations rather than relying solely on feature-space matching.
@@
Taken together, these findings indicate that Repa promotes perceptually meaningful reconstructions by steering the model toward regions of the feature space that reflect the semantic structure of natural images. Empirical evaluations in Section 5 reinforce this interpretation across multiple solvers and inverse problems.
@@ 5 Experimental Results
In this section, we present experimental results that demonstrate the effectiveness of our alignment regularizer. We evaluated reconstruction quality using four widely adopted metrics: PSNR (peak signal-to-noise ratio), SSIM (structure similarity index) (wang2004image), LPIPS (learned perceptual image patch similarity) (zhang2018unreasonable), and FID (Fréchet Inception Distance) (heusel2017gans). All metrics are averaged over 100 images from the ImageNet $256\times 256$ validation set (deng2009imagenet).
We consider four inverse problems: super-resolution, box inpainting, Gaussian deblurring, and motion deblurring. Following the standard corruption operators used in state-of-the-art diffusion-based inverse problem methods (chung2023diffusion; zhang2018unreasonable) we adopt the same degradation models in our experiments. For super-resolution, images are downsampled by a factor of $4$. For box inpainting, we mask out a $128\times 128$ square region. Gaussian deblurring is performed using a $61\times 61$ kernel with standard deviation $3.0$. Motion blur is simulated using randomly generated kernels with intensity $0.5$. In all experiments we also add Gaussian noise with standard deviation 0.01.
5.1 Effectiveness of the Repa Regularizer
This subsection evaluates how the Repa regularizer influences reconstruction quality by applying it to two representative latent-space solvers. To this end, we instantiate Algorithm 1 with Latent DPS and ReSample (song2024solving), which differ in how the measurement-matching term is approximated. A detailed description of these methods and their integration with Repa is provided in Appendix A.2. In all experiments, we use the latent diffusion model trained with representation alignment from yu2024repa, together with its pretrained MLP that projects internal diffusion representations into the DINOv2 feature space. Following yu2024repa, the regularizer is applied to the representations extracted after the eighth transformer block.
Table 1 reports quantitative results across all four inverse problems. Incorporating REPA yields consistent improvements in perceptual metrics notably reducing LPIPS and FID for both Latent DPS and ReSample while maintaining comparable PSNR and SSIM values. These results suggest that representation alignment guides the diffusion trajectory toward semantically consistent and perceptually realistic reconstructions that better match the target image. Figure 2 presents qualitative comparisons for box inpainting and Gaussian deblurring. Reconstructions obtained with Repa are visually sharper and more faithful to the ground truth, exhibiting fewer artifacts and clearer fine structures than those produced by the unregularized methods. Overall, these results confirm that applying Repa as an inference-time regularizer delivers both quantitative and perceptual gains, validating the proposed alignment mechanism as a general improvement for diffusion-based inverse problem solvers.
Table 1: Performance comparison of Latent DPS and ReSample (with and without the proposed Repa regularizer) on $4\times$ super-resolution, box inpainting, Gaussian deblurring, and motion deblurring tasks (ImageNet validation). Gaussian noise with $\sigma=0.01$ was added in all cases.
Task Method LPIPS ↓ FID ↓ PSNR ↑ SSIM ↑
$4\times$ Super-resolution Latent DPS 0.238 123.82 26.88 0.732
Latent DPS + REPA 0.217 86.88 26.82 0.731
Resample 0.208 97.02 26.60 0.732
Resample + REPA 0.197 74.70 26.41 0.724
Box Inpainting Latent DPS 0.151 116.53 20.53 0.827
Latent DPS + REPA 0.139 88.69 20.45 0.824
Gaussian Deblurring Latent DPS 0.288 152.96 25.76 0.648
Latent DPS + REPA 0.256 102.99 25.66 0.669
Resample 0.259 115.47 26.19 0.699
Resample + REPA 0.223 89.67 25.99 0.695
Motion Deblurring Latent DPS 0.249 129.08 27.19 0.738
Latent DPS + REPA 0.225 90.23 27.01 0.735
Resample 0.210 94.95 27.27 0.736
Resample + REPA 0.192 75.11 27.18 0.738
Measurement Latent–DPS + REPA Reference
@@ 5.1 Effectiveness of the Repa Regularizer
(a) Box inpainting
Measurement Latent–DPS + REPA Reference
@@
(b) Gaussian deblurring
Figure 2: Qualitative comparison of Latent DPS with and without the proposed REPA regularizer on two inverse problems: (a) box inpainting and (b) Gaussian deblurring.
@@
MeasurementDPSLatent DAPSReSample + REPAGround Truth
Figure 4: Qualitative comparison for motion deblurring. Each column shows the measurement, DPS, Latent DAPS, ReSample + REPA, and the ground-truth image.
Table 2: Comparison of our REPA-regularized reconstructions against state-of-the-art inverse problem solvers on four tasks ($4\times$ super-resolution, box inpainting, Gaussian deblurring, and motion deblurring) using ImageNet validation images. Gaussian noise with $\sigma=0.01$ is added in all cases. This table complements Table 1 by benchmarking REPA-enhanced models against external state-of-the-art baselines.
Task Method LPIPS ↓ FID ↓ PSNR ↑ SSIM ↑
$4\times$ Super-resolution DPS 0.244 92.13 24.42 0.681
Latent DAPS 0.252 120.57 27.13 0.744
Latent DPS + REPA 0.217 86.88 26.82 0.731
Resample + REPA 0.197 74.70 26.41 0.724
Box Inpainting DPS 0.191 97.95 19.11 0.769
Latent DPS + REPA 0.139 88.69 20.45 0.824
Gaussian Deblurring DPS 0.366 157.73 19.55 0.461
Latent DAPS 0.300 163.36 26.02 0.692
Latent DPS + REPA 0.256 102.99 25.66 0.669
Resample + REPA 0.223 89.67 25.99 0.695
Motion Deblurring DPS 0.242 89.06 24.17 0.678
Latent DAPS 0.274 135.18 27.01 0.734
Latent DPS + REPA 0.225 90.23 27.01 0.735
Resample + REPA 0.192 75.11 27.18 0.738
5.2 Comparison with other state of the art algorithms
After establishing that the REPA regularizer improves reconstruction quality, we compare our approach with other state-of-the-art algorithms. For pixel-space methods, we consider DPS (chung2023diffusion), which performs gradient-based sampling and is evaluated using the pretrained diffusion model from dhariwal2021diffusion. For latent diffusion approaches, we include Latent DAPS. Since our method relies on a model trained on ImageNet, we evaluated Latent DAPS with the standard conditional latent diffusion model trained on ImageNet from rombach2022high, to ensure a fair comparison. All aforementioned algorithms are run with their default configurations. Table 2 presents the quantitative results, where the inclusion of our regularizer consistently achieves state-of-the-art performance. Figure 4 provides qualitative examples for motion deblurring, further highlighting the advantage of our method. In particular, DPS often introduces visible artifacts, whereas Latent DAPS fails to fully correct blur and produces reconstructions lacking sharpness.

### v2

@@ 4 Theoretical Results
Recent advances in generative modeling have demonstrated that objectives defined in deep feature spaces, rather than purely pixel-based discrepancies, produce reconstructions that better capture perceptual similarity (Dosovitskiy and Brox, 2016; Zhang et al., 2018; Fu et al., 2023). Motivated by this observation, our goal in this section is to clarify how the Repa regularizer fits within this broader class of feature-based objectives. In particular, we present two theoretical results that characterize how Repa acts on (i) the DINOv2-space representations of reconstructed images and (ii) the latent diffusion state itself.
First, we show that the Repa term can be viewed as an empirical surrogate for minimizing a feature-space divergence specifically, the maximum mean discrepancy (MMD) between the DINOv2 features of reconstructed images and those of the ground truth. Second, we study how Repa affects the latent states of the diffusion model during the reverse process. Under suitable assumptions, we show that a Repa update moves the current diffusion state toward the latent representation of the clean image. This provides a complementary view: Repa influences not only the DINOv2 representation of the reconstruction but also the latent state that guides the sampling trajectory.
@@
Definition 4.1 (Misalignment between DINOv2 and diffusion representations).
@@
$\mathrm{MisREPA}(x,t):=\mathbb{E}_{n}\,\|f_{\!\mathrm{DINOv2}}^{[n]}(x)-g_{\phi}(\mathrm{DiffEnc}^{[n]}(x,t))\|_{2}^{2},$
where the expectation is taken uniformly over patch indices $n\in\{1,\dots,N\}$. Here $f_{\mathrm{DINOv2}}^{[n]}(x)$ and $\mathrm{DiffEnc}^{[n]}(x,t)$ denote the $n$-th patch features from the DINOv2 encoder and the diffusion model encoder, respectively. The mapping $g_{\phi}$ is the projection head introduced in Section 2.2111When the timestep argument is omitted, all $t$-dependent quantities are evaluated at $t=0$. In particular, $\mathrm{MisREPA}(x)\equiv\mathrm{MisREPA}(x,0)$ and $\mathrm{DiffEnc}^{[n]}(x)\equiv\mathrm{DiffEnc}^{[n]}(x,0)$..
Definition 4.2 (Proxy approximation error).
For a ground-truth image $x$ and its proxy approximation $\bar{x}$ (defined in Section 3), we define the proxy approximation error as
$\mathrm{ApproxErr}(x,\bar{x}):=\mathbb{E}_{n}\,\bigl\|f_{\mathrm{DINOv2}}^{[n]}(x)-f_{\mathrm{DINOv2}}^{[n]}(\bar{x})\bigr\|_{2}^{2},$
where the expectation is taken uniformly over patch indices $n\in\{1,\dots,N\}$.
In practice, both quantities defined above are typically small. The feature misalignment $\mathrm{MisREPA}(x,t)$ is mitigated through the explicit alignment of DINOv2 and diffusion encoder representations during training (Yu et al., 2024), while the approximation error $\mathrm{ApproxErr}(x,\bar{x})$ is reduced by constructing proxy features $\bar{x}$ that serves as reliable substitutes for the unavailable ground-truth features (see Section 3). Robustness of DINOv2 features to common image degradations (Oquab et al., 2023), ensures that the resulting approximation gap remains small across the inverse problems we study.
Feature-Space Alignment: Repa as an MMD Surrogate. Building on the definitions introduced above, we now establish a connection between the Repa regularizer and the statistical discrepancy between DINOv2 features of clean and reconstructed images. The proposition that follows formalizes this link by expressing the expected Repa score in terms of a feature-space divergence.
Proposition 4.3.
Assume $x\sim p_{X}$ is a ground-truth image and $\hat{x}\sim p_{\hat{X}\mid Y}$ is its corresponding reconstruction, where $p_{\hat{X}\mid Y}$ is the distribution induced by any reconstruction method. Assume further that all DINOv2 feature embeddings are $\ell_{2}$–normalized. Let $\bar{x}$ be a proxy approximation of $x$. Define the empirical mean DINOv2 feature embedding as
$\mu_{f}(x):=\mathbb{E}_{n}\!\left[f_{\mathrm{DINOv2}}^{[n]}(x)\right].$
Then the expected REPA alignment satisfies
@@
The proof is provided in the Appendix A.1. Proposition 4.3 establishes a connection between the representation alignment term and the maximum mean discrepancy (MMD) (Arbel et al., 2019), which quantifies perceptual divergence between feature distributions. This relationship indicates that maximizing the Repa term encourages the minimization of a divergence measure captured by the MMD together with additional penalties accounting for feature misalignment between diffusion and DINOv2 representations, as well as an approximation error arising from the use of proxy rather than ground-truth features for alignment.
@@
When these terms are sufficiently small, Proposition 4.3 implies that maximizing the Repa term effectively corresponds to minimizing the maximum mean discrepancy (MMD) between the DINOv2 feature distributions of reconstructed and ground-truth images. This interpretation reveals that Repa regularization offers a principled, representation-based mechanism for enforcing perceptual consistency.
Effect of REPA on Internal Diffusion Representations. We now study how a single Repa update influences the latent diffusion state relative to that of the clean image.
Proposition 4.4 (informal).
Consider an inverse problem with observation $y\in\mathbb{R}^{m}$ generated from a clean image $x\sim p_{X}$, and let $\bar{x}$ denote the proxy image used by REPA. Let $z_{t}\in\mathbb{R}^{d}$ be the diffusion state at time $t$, $z_{t}^{(\mathrm{REPA})}$ its REPA-regularized update. Let $z^{*}\in\mathbb{R}^{d}$ denote the latent variable associated with the clean image $x$. Under regularity assumptions on the Jacobian $J_{\mathrm{DiffEnc}}$ of the diffusion encoder $\mathrm{DiffEnc}$ and the mapping $g_{\phi}$ (see Appendix A.2), there exists a radius $r>0$ such that, whenever $\|z_{t}-z^{*}\|_{2}<r$, the REPA update induces a local contraction in the diffusion state space. In particular,
$\|z_{t}^{(\mathrm{REPA})}-z^{*}\|_{2}\leq C_{1}\,\|z_{t}-z^{*}\|_{2}+e(x),$
$e(x)$ $=C_{2}\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right),$
where $C_{1}\in(0,1)$ and $C_{2}>0$ are constants depending on local conditioning properties of $J_{\mathrm{DiffEnc}}$ and $g_{\phi}$.
The proof is provided in Appendix A.1. Proposition 4.4 characterizes how the Repa update modifies the latent diffusion state. The inequality shows that, when the current state $z_{t}$ is reasonably close to the clean latent code $z^{*}$, a single Repa step contracts their distance up to two additive terms: the proxy approximation error $\mathrm{ApproxErr}(x,\bar{x})$ and the representation misalignment $\mathrm{MisREPA}(x)$, both of which are expected to be small in practice. This contraction therefore holds in a local regime around $z^{*}$, which corresponds in practice to stages of the diffusion process where the global structure of the reconstruction is already correct and the remaining errors are primarily fine-scale details and artifacts. In this regime, Repa acts as a perceptual regularizer, refining local image details and correcting artifacts that are not effectively addressed by the baseline update alone. Qualitative examples illustrating this behavior are provided in Figure 2 and the Appendix A.9.
@@
Taken together, Propositions 4.3 and 4.4 indicate that Repa promotes perceptually meaningful reconstructions by steering the model toward regions of the latent space that reflect the semantic structure of natural images. Empirical evaluations in Section 5 reinforce this interpretation across multiple solvers and inverse problems.
@@ 5 Experimental Results
Measurement Without REPA With REPA Reference
@@
Measurement Without REPA With REPA Reference
@@
Measurement Without REPA With REPA Reference
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Measurement Without REPA With REPA Reference
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Qualitative comparison of inverse problem reconstruction with and without Repa. The top block corresponds to box inpainting and the bottom block to Gaussian deblurring; left panels show ImageNet results and right panels show FFHQ results. For each example, we show the measurement, the baseline reconstruction (Latent DPS or ReSample), the corresponding Repa-enhanced result, and the reference image. Within each panel, the first row corresponds to Latent DPS and the second row to ReSample.
In this section, we present experimental results demonstrating the effectiveness of our alignment regularizer. We evaluate reconstruction quality using four widely adopted metrics: PSNR (peak signal-to-noise ratio), SSIM (structural similarity index) (Wang et al., 2004), LPIPS (learned perceptual image patch similarity) (Zhang et al., 2018), and FID (Fréchet Inception Distance) (Heusel et al., 2017). Experiments are conducted on both the ImageNet and FFHQ datasets (Deng et al., 2009; Karras et al., 2019), with FFHQ images resized to $256\times 256$. All metrics for both datasets are averaged over 100 images from their corresponding validation splits.
We consider four inverse problems: super-resolution, box inpainting, Gaussian deblurring, and motion deblurring. Following the standard corruption operators used in state-of-the-art diffusion-based inverse problem methods (Chung et al., 2023; Zhang et al., 2025) we adopt the same degradation models in our experiments. For super-resolution, images are downsampled by a factor of $4$. For box inpainting, we mask out a $128\times 128$ square region. Gaussian deblurring is performed using a $61\times 61$ kernel with standard deviation $3.0$. Motion blur is simulated using a kernel with size $61\times 61$ and intensity $0.5$. For ImageNet we use additive Gaussian noise with standard deviation $0.01$. For FFHQ we increase the noise level to $0.05$ in order to obtain a more challenging reconstruction setting.
Table 1: Performance comparison on ImageNet and FFHQ across inverse tasks. ImageNet experiments use Gaussian noise with $\sigma=0.01$, while FFHQ uses $\sigma=0.05$. Lower is better for LPIPS/FID, higher is better for PSNR/SSIM. Best per dataset and task in bold.
ImageNet FFHQ
Task Method LPIPS↓ FID↓ PSNR↑ SSIM↑ LPIPS↓ FID↓ PSNR↑ SSIM↑
$4\times$ SR Latent DPS 0.238 123.82 26.88 0.732 0.188 56.69 28.99 0.814
Latent DPS + REPA 0.217 86.88 26.82 0.731 0.177 51.27 29.12 0.819
Resample 0.208 97.02 26.70 0.736 0.178 52.82 29.06 0.814
Resample + REPA 0.197 74.70 26.99 0.740 0.167 46.36 29.17 0.818
DPS 0.244 92.13 24.42 0.681 0.182 54.19 27.67 0.804
Latent DAPS 0.252 120.57 27.13 0.744 0.206 74.82 28.95 0.820
Inpainting Latent DPS 0.151 116.53 20.53 0.822 0.192 65.89 23.55 0.785
Latent DPS + REPA 0.139 88.69 20.45 0.824 0.178 57.04 23.83 0.784
Resample 0.153 121.98 20.53 0.812 0.192 67.99 23.86 0.792
Resample + REPA 0.143 87.15 20.79 0.827 0.178 61.97 24.01 0.793
DPS 0.191 97.95 19.11 0.769 0.190 104.53 20.02 0.786
Latent DAPS 0.318 188.44 21.21 0.718 0.210 89.21 24.39 0.817
Gaussian Deblur Latent DPS 0.288 152.96 25.76 0.648 0.192 60.65 28.15 0.783
Latent DPS + REPA 0.256 102.99 25.66 0.669 0.186 55.53 28.21 0.787
Resample 0.259 115.47 26.19 0.699 0.172 56.41 27.01 0.753
Resample + REPA 0.223 89.67 26.49 0.707 0.168 52.73 27.17 0.756
DPS 0.366 157.73 19.55 0.461 0.177 53.18 27.19 0.789
Latent DAPS 0.291 159.67 26.02 0.692 0.229 104.43 28.73 0.809
Motion Deblur Latent DPS 0.249 129.08 27.19 0.738 0.170 52.14 27.20 0.773
Latent DPS + REPA 0.225 90.23 27.01 0.735 0.165 47.18 27.16 0.772
Resample 0.210 89.95 27.57 0.739 0.157 52.19 28.36 0.791
Resample + REPA 0.192 75.11 27.80 0.766 0.151 50.02 28.41 0.794
DPS 0.242 89.06 24.17 0.678 0.146 43.32 27.28 0.793
Latent DAPS 0.264 125.18 27.25 0.744 0.191 76.21 29.77 0.836
@@
5.1 Effectiveness of the Repa Regularizer
This subsection evaluates how the Repa regularizer influences reconstruction quality by applying it to two representative latent-space solvers. To this end, we instantiate Algorithm 1 with Latent DPS and ReSample (Song et al., 2024), which differ in how the measurement-matching term is approximated. A detailed description of these methods and their integration with Repa is provided in Appendix A.2. For ImageNet experiments, we use the latent diffusion model trained with representation alignment from Yu et al. (2024), together with its pretrained MLP that projects internal diffusion representations into the DINOv2 feature space. For FFHQ, we train a representation-aligned SiT-Base model from scratch following the same setup as Yu et al. (2024). In ImageNet experiments, the regularizer is applied to representations extracted after the eighth transformer block, while for FFHQ it is applied after the fourth block. For completeness, Appendix A.8 additionally reports results obtained when starting from an independently pretrained diffusion model. In this setting, we observe that incorporating the Repa regularizer continues to provide consistent perceptual benefits, although the gains are smaller compared to those achieved with alignment-pretrained models
Table 1 reports quantitative results across all four inverse problems. Incorporating Repa yields consistent improvements in perceptual metrics, notably reducing LPIPS and FID for both Latent DPS and ReSample while maintaining comparable PSNR and SSIM values. These results suggest that representation alignment guides the diffusion trajectory toward semantically consistent and perceptually realistic reconstructions that better match the target image. Figure 2 presents qualitative comparisons for box inpainting and Gaussian deblurring, where reconstructions obtained with Repa are visually sharper and more faithful to the ground truth, exhibiting fewer artifacts and clearer fine structures than those produced by the unregularized methods.
In addition to improving these latent solvers, we also compare our approach with other state-of-the-art reconstruction algorithms. For pixel-space methods, we consider DPS (Chung et al., 2023), evaluated using the pretrained diffusion model from Dhariwal and Nichol (2021). For latent diffusion baselines, we include Latent DAPS, implemented with the conditional ImageNet latent diffusion model from Rombach et al. (2022). As shown in Table 1, the proposed alignment regularizer consistently achieves state-of-the-art performance in perception-based metrics across all inverse tasks and both datasets.
Repa vs Guidance in the DINO feature space. To further showcase the benefits of representation alignment, Appendix A.7 compares Repa to a variant that enforces consistency directly in the DINOv2 feature space of the reconstructed image, without aligning the diffusion model’s internal states. Concretely, the feature-only variant applies the following guidance at each timestep:
$z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\Bigl(c^{[n]}_{\text{proxy}},f_{\mathrm{DINOv2}}^{[n]}\!\bigl(\hat{x}_{0}(z_{t})\bigr)\Bigr),$
where $\hat{x}_{0}(z_{t}):=\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])$ denotes the decoded prediction of the clean image at timestep $t$. This approach relies solely on the DINOv2 features of the decoded image. In contrast, Repa injects representation alignment directly into the diffusion dynamics by aligning the model’s internal representations with DINOv2 features throughout the reverse process. As observed in Appendix A.7, this feature-only guidance performs worse in practice. A natural explanation trough the lens of Proposition 4.4 is that the composite mapping $z_{t}\mapsto\mathbb{E}[z_{0}\mid z_{t}]\mapsto\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\mapsto f_{\mathrm{DINOv2}}(\cdot)$ is significantly more ill-conditioned than the lightweight patchwise projection used in Repa, resulting in weaker and less stable gradients with respect to the diffusion state.